one day at planet X is 36 hours. The mass of planet X is 5.55.10^24 kg.

Part A: Find radius of orbit of a geosynchronous satellite that circles planet X ?
Answer #1

Using Kepler's 3rd law:

T^2 = 4*pi^2*R^3/(G*M)

We know that, M = mass of planet = 5.55*10^24 kg

G = 6.67*10^-11

T = time period of geosynchronous satellite = 36 hrs = 129600 sec

So,

R = [GMT^2/(4*pi^2)]^(1/3)

R = [6.67*10^-11*5.55*10^24*129600^2/(4*pi^2)]^(1/3)

R = 5.40*10^7 m = radius of orbit of geo satellite